# Seeing Through Green: Text-Based Classification and the Firm’s Returns from Green Patents

### v1

@@
Seeing Through Green: Text-Based Classification and the Firm’s Returns from Green Patents
Lapo Santarlasci Mail to: lapo.santarlasci@imtlucca.it. Laboratory for the Analysis of Complex Economic Systems (AXES), IMT School for Advanced Studies, Lucca    Armando Rungi Mail to: armando.rungi@imtlucca.it. Laboratory for the Analysis of Complex Economic Systems (AXES), IMT School for Advanced Studies, Lucca    Antonio Zinilli Mail to: antonio.zinilli@cnr.it. National Research Council of Italy (CNR), Research Institute on Sustainable Economic Growth, Unit of Rome
This paper introduces Natural Language Processing for identifying ’true’ green patents from official supporting documents. We start our training on about 12.4 million patents that had been classified as green from previous literature. Thus, we train a simple neural network to enlarge a baseline dictionary through vector representations of expressions related to environmental technologies. After testing, we find that ’true’ green patents represent about 20% of the total of patents classified as green from previous literature. We show heterogeneity by technological classes, and then check that ’true’ green patents are about 1% less cited by following inventions. In the second part of the paper, we test the relationship between patenting and a dashboard of firm-level financial accounts in the European Union. After controlling for reverse causality, we show that holding at least one ’true’ green patent raises sales, market shares, and productivity. If we restrict the analysis to high-novelty ’true’ green patents, we find that they also yield higher profits. Our findings underscore the importance of using text analyses to gauge finer-grained patent classifications that are useful for policymaking in different domains.
@@ 1 Introduction
Climate change is no longer a distant externality; it is a binding economic constraint that is already reshaping product markets, capital allocation and regulatory priorities worldwide. In the European Union, for example, the Green Deal, the Fit-for-55 package and a growing array of national carbon-pricing schemes collectively tighten the link between a firm’s environmental footprint and its license to operate. Against this backdrop, green innovation has emerged as a central lever for reconciling competitiveness with decarbonisation goals. Yet a decade of empirical work has produced mixed verdicts on whether green innovation translates into superior firm performance. While several studies document positive stock-market reactions or profitability gains, others find negligible or even negative effects once R&D crowd-out and compliance costs or demand uncertainty are considered (Barbieri et al., 2017; Przychodzen and Przychodzen, 2015).
@@
In this contribution, we start by challenging standard classifications of green patents, i.e., widely used lists such as OECD’s ENV-TECH, the EPO’s Y02/Y04S and WIPO’s IPC Green Inventory assign green status to entire patent subclasses. We believe that previous approaches guarantee coverage but risk conflating genuinely environmentally efficient inventions with conventional technologies that merely happen to sit in an environmentally relevant class. In this paper, we propose the adoption of a Natural Language Processing method based on a simple neural network algorithm that allows us to end up with a dictionary for environmental technologies. Through a so-called continuous bag of words, the algorithm uses a shallow neural network with a single hidden layer. The model learn semantic relationships from patents titles and abstracts, and leverage them to enlarge the seed dictionary derived from large language models, with synonyms. The model computes the 15 most semantically similar expressions for each seed. For our purpose, we train our algorithm on a unique database of about 12,400,000 patents. The result is a selection of what we consider true green patents, which is substantially narrower, i.e., 20%, than the one based on existing classifications.
@@
We believe our proposal is particularly useful in combination with another text-based indicator proposed by Arts et al. (2021), which catches the novelty of patents based on previously unseen keywords and phrases. What we want to avoid with this implementation is for greenwashing to hide in plain sight. As regulatory and investor attention intensifies, firms may strategically patent marginal design tweaks to signal sustainability credentials — a form of greenwashing that inflates the counts of green patents without a corresponding technological progress in environmental protection. Once we consider also the novelty degree of true green patents, we find that they constitute about 49% of the total.
@@
We find that holding at least one green patent is associated with more than double sales, a market share that is higher by 38%, and productivity gains by 40%. If we further check for the novelty degree of the green patents, after we interact with the indicator by Arts et al. (2021), we find that high-novelty green patents have almost double the gains on sales. Previous results can be due to a combination of protection of property rights and the innovative content of the invention, both of which provide an explanation for a positive association with firms’ financial accounts (Exadaktylos et al., 2024).
The remainder of the paper is structured as follow: in Section 2 we present the related literature highlighting our contribution and the reason behind our research questions; in Section 3 we present the dataset we constructed and we show some descriptive evidence; in Section 4 we the present our custom NLP classification procedures and its results; in Section 5 we present our empirical strategy and final results, and, finally, in Section 6 some robustness and sensitivity checks.
@@ 2 Literature review
Patents are widely viewed as a key output of innovation and can significantly contribute to economic value creation (Artz et al., 2010; Howell, 2017). Nonetheless, patent quality and impact vary considerably; while most patents represent incremental advancements, only a small subset introduces radical innovation (Artz et al., 2010), and empirical evidence points to a positive correlation between higher-quality patents and stronger financial or economic performance (Balasubramanian and Sivadasan, 2011; Hirshleifer et al., 2013; Kogan et al., 2017). To assess this quality, researchers often use patent citations as an indicator of both technological and economic value (Trajtenberg, 1990; Hall et al., 2005; Kogan et al., 2017; Andriosopoulos et al., 2022). Additional measures include patent scope (Barbieri et al., 2017), patent family size (Lanjouw and Mody, 1996; Hall and Helmers, 2013), and the count of non-patent references (Callaert et al., 2006; Hall and Helmers, 2013). Of particular interest is the work by (Arts et al., 2021), which introduces NLP-based measures that often outperform traditional classification and citation-based approaches. These new measures, for example, more accurately identify patents associated with prestigious awards (indicative of high impact) and patents granted by the USPTO but rejected by the EPO and JPO (suggestive of low novelty). Specifically, we believe that it is important to correctly measure green patents and their novelty to avoid greenwashing [^2] Greenwashing refers to “misleading the public to believe that a company or entity is doing more to protect the environment than it is” (United Nations, 2022) practices. Firms may engage in greenwashing to bolster their environmental legitimacy and gain support from consumers and green investors (Berrone et al., 2010, 2017), or they might seek to cultivate a green image (Cai and Li, 2018). While green patents can signal environmental progress to investors (Amore and Bennedsen, 2016), their market valuation depends on firm characteristics and on the specific attributes of the patent (Chirico et al., 2025).
@@ 3.1 Patent data
To initiate the classification process, we employ three broadly acknowledged algorithms to source patents from PATSTAT: the ENV-TECH, developed by the Organisation for Economic Co-operation and Development (OECD, Haščič and Migotto (2015)), the EPO “Y02/Y04S tagging scheme” (Angelucci et al., 2018) and the WIPO and UNFCCC “IPC Green Inventory” [^3] https://www.wipo.int/classifications/ipc/green-inventory/home. They operate based on the International Patent Classification (IPC) or the Cooperative Patent Classification (CPC), which consider patents as attached to specific technological fields. Table 1 shows an example of IPC codes that the ENV-TECH algorithm links to green patents.
@@
Although developed independently, the three methodologies are complementary since they cover non perfectly overlapping technological domains, so we use them in combination to obtain a wider coverage for our exercise. Table 2 summarizes the main differences between these methodologies.
@@ 3.2 Firm-level data
In the second part of our paper, we reduce the geographical scope of our analysis and study how green patenting is associated with firms’ competitiveness in the European Union. We select all companies with available data to investigate an entire dashboard of firms’ financial accounts in the period 2010-2022. At the beginning, we consider all firms, whether they have patents or not. We end up with a sample of 2,676,515 European firms. Appendix Table B1 shows the number of companies by country, while Appendix Table B2 shows the classification of firms by number of employees. Notably, in Appendix Tables B3, we show that our firm-level match covers all 2-digit industries. When we select only those that have registered patents in our period of analysis, we have a subsample of 15,462 companies. Importantly, among them, only 3,700 have one green patent according to previous classification in the literature, and if we look at the set of firms that have at least one true green patent in the same period, we end up with only 796 firms. For our purpose, we design a dashboard of firms’ outcomes that allows us to understand the gains from green patenting. For each firm, when available, we select sales, market share, labor productivity, capital intensity, ROCE (Return on Capital Employed), EBIT (Earnings Before Interest and Taxes), and TFP (Total Factor Productivity).
@@ 4 A Natural Language Processing approach to classify true green patents
Although classification codes are widely used for separating green patents, they can lead to cases of misclassification. We report anecdotal evidence in Table 3. The code F01N7 “Exhaust or silencing apparatus, or parts thereof” from the ENV-TECH list can surely be associated with a technical field in which innovation often leads to less polluting engines, as a result of the use of more advanced exhaust systems. But the same code can include an invention for a component that increases engine performance by burning a larger amount of fuel, regardless of environmental concerns.
Table 3: Example IPC classes used by the ENV-TECH algorithm
IPC Description
F01N7 Exhaust or silencing apparatus, or parts thereof
F02M39-71 Fuel-injection apparatus
  • Note: The table reports examples of controversial IPC classes used by the ENV-TECH algorithm for the classification of green patents.
@@
Our custom vocabulary for environmental technologies is finally used to perform regex matches on patent descriptions, on the sub-sample of patents classified as green by the code-based algorithms, enabling us to identify patents that are more closely aligned with the concept of green innovation. We apply it to abstracts and titles of patents in English [^7] A few patents are not written in English, and we translate them with the help of the Python translator, which systematically leverages Google Translate, after translation, we also adopt NLP techniques to: remove stop words, and apply part-of-speech tagging to retain only nouns, adjectives, verbs, adverbs and proper nouns. We convert to lowercase, lemmatize every word, replace numbers and measures (e.g., LxW or LxWxH) with a generic tag, and remove special characters. Finally, we concatenate titles and abstracts into unique texts
@@ 4.1.1 Share of true greens in technological classes
The final result of our NLP approach is a vocabulary tailored to the domain of green technologies. In Appendix Tables A1 A2, A3 and A4, we report a comparison between the number of green patents obtained from previous classifications in each CPC code and the number of true green patents that we obtain with our approach. In Figure 2, we summarize the shares of true greens for each CPC code and their density. On average, we observe a 20% share of the total, but with a dispersion that runs from 0 to 100% [^8] It is important to highlight that a single invention effort can be acknowledged as a patent in more than one technological class, hence descriptive statistics here include multiple counting..
When shifting to a more granular level of classification, we notice how, specifically, CPCs related to “Technologies or applications for mitigation or adaptation against climate change” (Y02), “Electric elements” (H01) and “Physical or chemical processes or apparatus in general” (B01) are the ones with the highest number of green patents if we believe in previous classifications. When we use our approach, CPC Y02 remains the category associated with the largest number of true greens, but then we find “Generation, conversion, or distribution of electric power” (H02) and “Treatment of water, waste water, sewage, or sludge” (C02) in the second and third positions. If we look instead at the share of true green patents we see “Sugar industry” (C13), “Doors, windows, shutters, or roller blinds, in general; ladders” (E06) and “Heating; ranges; ventilating” (F24) in the top 3 rank with shares between 40% and 50%.
@@
  • Note: The figure represents the density of the shares of true green patents vs the total of green patents identified by previous classifications.
4.1.2 Technological classes and their specialisation in green patents
@@
We computed the RCA index for each CPC/IPC category at a 1- and 3-digit level. Figure 3 shows the density distribution of the RCA index values by technological domains. While the distribution is centred below 1, indicating that green patents are on average more present in many non-specialised categories, where the left tail is slightly left-skewed. In other words, in general, there are green patents across many technological domains, whose main scope, however, is not to produce environmental technologies. When shifting to a more granular level of classification (please refer to Appendix B.2. Classification results, Tables B9, B10 and B11 for complete numbers of the RCA index at 3-digit of the CPC code), we observe how, specifically, top 3 positions are occupied by CPC Treatment of water, waste water, sewage, or sludge (C02), Technologies or applications for mitigation or adaptation against climate change (Y02), and Disposal of solid waste; reclamation of contaminated soil (B09).

### v2

@@
Seeing Through Green: Text-Based Classification and the Firm’s Returns from Green Patents thanks: Authors claim funding from the European Union - NEXT-GENERATION EU, as part of the project “Fostering Open Science in Social Science Research (FOSSR)” - MUR Code PNRR-IR0000008 - CUP B83C22003950001. We would like to thank Maurizio Zanardi for his insightful suggestions.
Lapo Santarlasci Mail to: lapo.santarlasci@imtlucca.it. Laboratory for the Analysis of Complex Economic Systems (AXES), IMT School for Advanced Studies, Lucca    Armando Rungi Mail to: armando.rungi@imtlucca.it. Laboratory for the Analysis of Complex Economic Systems (AXES), IMT School for Advanced Studies, Lucca    Antonio Zinilli Mail to: antonio.zinilli@cnr.it. National Research Council of Italy (CNR), Research Institute on Sustainable Economic Growth, Unit of Rome.
This paper introduces Natural Language Processing for identifying ’true’ green patents from official supporting documents. We start our training on about 12.4 million patents that had been classified as green from previous literature. Thus, we train a simple neural network to enlarge a baseline dictionary through vector representations of expressions related to environmental technologies. After testing, we find that ’true’ green patents represent about 18.5% of the total of patents classified as green from previous literature. We show heterogeneity by technological classes, and then check that ’true’ green patents are about 4% less cited by following inventions. In the second part of the paper, we test the relationship between patenting and a dashboard of firm-level financial accounts in the European Union. After controlling for reverse causality, we show that holding at least one ’true’ green patent raises sales, market shares, and productivity. If we restrict the analysis to high‑novelty ’true’ green patents, we find that they also yield higher profits. Our findings underscore the importance of using text analyses to gauge finer‑grained patent classifications that are useful for policymaking in different domains.
@@ 1 Introduction
Climate change is no longer a distant externality; it is a binding economic constraint that is already reshaping product markets, capital allocation and regulatory priorities worldwide. In the European Union, for example, the Green Deal, the Fit‑for‑55 package and a growing array of national carbon‑pricing schemes collectively tighten the link between a firm’s environmental footprint and its license to operate. Against this backdrop, green innovation has emerged as a central lever for reconciling competitiveness with decarbonisation goals. Yet a decade of empirical work has produced mixed verdicts on whether green innovation translates into superior firm performance. While several studies document positive stock‑market reactions or profitability gains, others find negligible or even negative effects once R&D crowd‑out and compliance costs or demand uncertainty are considered (Barbieri et al., 2017; Przychodzen and Przychodzen, 2015).
@@
In this contribution, we start by challenging standard classifications of green patents, i.e., widely used lists such as OECD’s ENV-TECH and WIPO’s IPC Green Inventory assign green status to entire patent subclasses, or the EPO’s Y02/Y04S tagging scheme. While the latter also apply automated text-based algorithms for green patent classification, their methodology is proprietary and non-transparent: the specific keywords, search terms, and algorithmic structures are not publicly disclosed. This creates a transparency and reproducibility problem for academic research. We believe that previous approaches guarantee coverage but risk conflating genuinely environmentally efficient inventions with conventional technologies that merely happen to sit in an environmentally relevant class and/or do not provide transparent reproducible classification frameworks. In this paper, we propose the adoption of a Natural Language Processing method based on a simple neural network algorithm that allows us to end up with a dictionary for environmental technologies. Through a so-called continuous bag of words, the algorithm uses a shallow neural network with a single hidden layer. The model learn semantic relationships from patents titles and abstracts, and leverage them to enlarge the seed dictionary derived from large language models, with synonyms. The model computes the 15 most semantically similar expressions for each seed. For our purpose, we train our algorithm on a unique database of about 12,400,000 patents. The result is a selection of what we consider true green patents, which is substantially narrower, i.e., 18.5%, than the one based on existing classifications.
In addition to find that standard green patent classifications may suffer from Type I errors (false positives) when conventional technologies are misclassified due to broad technological class assignments, we identify Type II errors (false negatives) when genuine green innovations fall outside recognized environmental domains ($\sim 2.9\%$ during our sample period). This type of under-classification is inherently due to a limited or biased selection of certain technological domains by experts defining standard classification methods. Our NLP approach addresses these systematic biases by analyzing patent content directly, providing a more accurate foundation for understanding environmental innovation.
@@
We believe our proposal is particularly useful in combination with another text‑based indicator proposed by Arts et al. (2021), which catches the novelty of patents based on previously unseen keywords and phrases. What we want to avoid with this implementation is for greenwashing to hide in plain sight. As regulatory and investor attention intensifies, firms may strategically patent marginal design tweaks to signal sustainability credentials — a form of greenwashing that inflates the counts of green patents without a corresponding technological progress in environmental protection. Once we consider also the novelty degree of true green patents, we find that they constitute about 49% of the total.
@@
We find that holding at least one green patent is associated with more than double sales, a market share that is higher by 38%, and productivity gains by 41%. If we further check for the novelty degree of the green patents, after we interact with the indicator by Arts et al. (2021), we find that high‑novelty green patents have almost double the gains in sales. Previous results can be due to a combination of protection of property rights and the innovative content of the invention, both of which provide an explanation for a positive association with firms’ financial accounts (Exadaktylos et al., 2024).
The remainder of the paper is structured as follow: in Section 2 we present the related literature highlighting our contribution and the reason behind our research questions; in Section 3 we present the dataset we constructed and we show some descriptive evidence; in Section 4 we then present our custom NLP classification procedures and its results; in Section 5 we present our empirical strategy and final results, and, finally, in Section 6 some robustness and sensitivity checks.
@@ 2 Literature review
Patents are widely viewed as a key output of innovation and can significantly contribute to economic value creation (Artz et al., 2010; Howell, 2017). Nonetheless, patent quality and impact vary considerably; while most patents represent incremental advancements, only a small subset introduces radical innovation (Artz et al., 2010), and empirical evidence points to a positive correlation between higher-quality patents and stronger financial or economic performance (Balasubramanian and Sivadasan, 2011; Hirshleifer et al., 2013; Kogan et al., 2017). To assess this quality, researchers often use patent citations as an indicator of both technological and economic value (Trajtenberg, 1990; Hall et al., 2005; Kogan et al., 2017). Additional measures include patent scope (Barbieri et al., 2017), patent family size (Lanjouw and Mody, 1996; Hall and Helmers, 2013), and the count of non-patent references (Callaert et al., 2006; Hall and Helmers, 2013). Of particular interest is the work by (Arts et al., 2021), which introduces NLP-based measures that often outperform traditional classification and citation-based approaches. These new measures, for example, more accurately identify patents associated with prestigious awards (indicative of high impact) and patents granted by the USPTO but rejected by the EPO and JPO (suggestive of low novelty).
Specifically, we believe that it is important to correctly measure green patents and their novelty to avoid greenwashing [^2] Greenwashing refers to “misleading the public to believe that a company or entity is doing more to protect the environment than it is” (United Nations, 2022) practices. Firms may engage in greenwashing to bolster their environmental legitimacy and gain support from consumers and green investors (Berrone et al., 2010, 2017), or they might seek to cultivate a green image (Cai and Li, 2018). While green patents can signal environmental progress to investors (Amore and Bennedsen, 2016), their market valuation depends on firm characteristics and on the specific attributes of the patent (Chirico et al., 2025).
@@ 3.1 Patent data
To initiate the classification process, we employ three broadly acknowledged algorithms to source patents from PATSTAT: the ENV-TECH, developed by the Organisation for Economic Co-operation and Development (OECD, Haščič and Migotto (2015)), the EPO “Y02/Y04S tagging scheme” (Veefkind et al., 2012) and the WIPO and UNFCCC “IPC Green Inventory” [^3] https://www.wipo.int/classifications/ipc/green-inventory/home. The first two operate based on the International Patent Classification (IPC) or the Cooperative Patent Classification (CPC), which consider patents as attached to specific technological fields. Table 1 shows an example of IPC codes that the ENV-TECH algorithm links to green patents.
@@
While the EPO Y02/Y04S tagging scheme also leverage automated text-based algorithms for green patent classification, their methodology is proprietary and non-transparent. In fact, the EPO combines classification codes with expert-designed search algorithms, but the specific keywords, search terms, and algorithmic structures are not publicly disclosed. This creates a transparency and reproducibility problem for academic research. Although developed independently, the three methodologies are complementary since they cover non perfectly overlapping technological domains, so we use them in combination to obtain a wider coverage for our exercise. Table 2 summarizes the main differences between these methodologies.
@@ 3.2 Firm-level data
In the second part of our paper, we reduce the geographical scope of our analysis and study how green patenting is associated with firms’ competitiveness in the European Union. We select all companies with available data to investigate an entire dashboard of firms’ financial accounts in the period 2010-2022. At the beginning, we consider all firms, whether they have patents or not. We end up with a sample of 2,676,515 European firms. Appendix Table B1 shows the number of companies by country, while Appendix Table B2 shows the classification of firms by number of employees. Notably, in Appendix Tables B3, we show that our firm-level match covers all 2-digit industries. When we select only those that have registered patents in our period of analysis, we have a subsample of 15,462 companies. Importantly, among them, only 3,700 have one green patent according to previous classification in the literature, and if we look at the set of firms that have at least one true green patent in the same period, we end up with only 596 firms. For our purpose, we design a dashboard of firms’ outcomes that allows us to understand the gains from green patenting. For each firm, when available, we select sales, market share, labor productivity, capital intensity, ROCE (Return on Capital Employed), EBIT (Earnings Before Interest and Taxes), and TFP (Total Factor Productivity).
@@ 4 A Natural Language Processing approach to classify true green patents
Although classification codes are widely used for separating green patents, they can lead to cases of misclassification. For example, the code F01N7 “Exhaust or silencing apparatus, or parts thereof” from the ENV-TECH list can surely be associated with a technical field in which innovation often leads to less polluting engines, as a result of the use of more advanced exhaust systems. But the same code can include an invention for a component that increases engine performance by burning a larger amount of fuel, regardless of environmental concerns.
@@
Our custom vocabulary for environmental technologies ( Appendix Appendix A: Natural language processing Tables A1 A2, and A3) is finally used to perform regex matches on patent descriptions, on the sub-sample of patents classified as green by the standard classification algorithms, enabling us to identify patents that are more closely aligned with the concept of green innovation. We apply it to abstracts and titles of patents in English [^7] A few patents are not written in English, and we translate them with the help of the Python translator, which systematically leverages Google Translate, after translation, we also adopt NLP techniques to: remove stop words, and apply part-of-speech tagging to retain only nouns, adjectives, verbs, adverbs and proper nouns. We convert to lowercase, lemmatise every word, replace numbers and measures (e.g., LxW or LxWxH) with a generic tag, and remove special characters. Finally, we concatenate titles and abstracts into unique texts
@@ 4.1.1 Share of true greens in technological classes
In Appendix Appendix B: Tables and graphs, Tables B6, B7, and B8, we report a comparison between the number of green patents obtained from previous classifications in each CPC code and the number of true green patents that we obtain with our approach. In Figure 2, we summarize the shares of true greens for each CPC code and their density. On average, we observe a 18% share of the total, but with a dispersion that runs from 0 to 100% [^8] It is important to highlight that a single invention effort can be acknowledged as a patent in more than one technological class, hence descriptive statistics here include multiple counting..
When shifting to a more granular level of classification, we notice how, specifically, CPCs related to “Technologies or applications for mitigation or adaptation against climate change” (Y02 - from EPO Y-tagging scheme), “Generation, conversion, or distribution of electric power” (H02) and “Physical or chemical processes or apparatus in general” (B01) are the ones with the highest number of green patents if we believe in previous classifications. When we use our approach, the EPO tagging scheme derived CPC Y02 remains the category associated with the largest number of true greens, but then we find “Generation, conversion, or distribution of electric power” (H02) and “Treatment of water, waste water, sewage, or sludge” (C02) in the second and third positions. If we look instead at the share of true green patents we see “Sugar industry” (C13), “Heating; ranges; ventilating” (F24), and “Generation, conversion, or distribution of electric power” (H02) in the top 3 rank with shares between 35% and 50%.
@@
  • Note: The figure represents the density of the shares of true green patents by CPC category vs. the total of green patents identified by previous classifications.
4.1.2 Technological classes and their specialization in green patents
@@
We computed the RCA index for each CPC/IPC category at a 1- and 3-digit level. Figure 3 shows the density distribution of the RCA index values by technological domains. While the distribution is centred below 1, indicating that green patents are on average more present in many non-specialised categories, where the left tail is slightly left-skewed. In other words, in general, there are green patents across many technological domains, whose main scope, however, is not to produce environmental technologies. When shifting to a more granular level of classification (please refer to Appendix B.2. Classification results, Tables B9, B10 and B11 for complete numbers of the RCA index at 3-digit of the CPC code), we observe how, specifically, top 3 positions are occupied by CPC “Treatment of water, waste water, sewage, or sludge” (C02), “Disposal of solid waste; reclamation of contaminated soil” (B09), and “Technologies or applications for mitigation or adaptation against climate change” (Y02).
@@ 4.1.3 The dynamics of green patents
4.2 Matching outside green perimeter
We aim to identify if our methodology flags “green patents” that are missed by standard classification algorithms developed by major institutions. This check is crucial for uncovering another possible drawbacks of classical classifiers: underclassification. Our findings revealed over 1,026,000 granted applications (around 2.9% of the total non-green patents in the sample period) that our methodology identified as green, despite not being flagged by other approaches. As an example an application with CPC codes B08B3/02, B08B3/14, and B08B13/00 (related to cleaning technologies) includes an abstract stating: “Waste liquid recycling device for surface treatment cleaning … According to the wastewater treatment device, the liquid storage tank, the sealing gasket and the floating block are matched with one another, so that wastewater can be purified, the wastewater can be reutilized, and the loss is reduced.”. Another application, classified under A01H1/02, A01G7/06, A01G22/00, A01H1/027, and A01K49/00 (related to methods to promote plant growth), is described as: “Eco-friendly pollinator method for small scale crops - purpose: A method for pollination in small scale seed gathering place is provided to reduce labor cost and to collect seed… An environmentally-friendly pollination of small scale crops is performed additionally using male bees…”.
These examples suggest that a purely standard classification approach may lead to noisy classifications, potentially including non-green patents while excluding legitimate ones from less common technological domains. From a classification error perspective, our results suggest that traditional classification approaches suffer from substantial Type I errors (false positives), incorrectly labeling non-environmental patents as green, as evidenced by only 18.5% of traditionally classified green patents meeting our “true green” criteria. Conversely, our NLP methodology appears to address significant Type II errors (false negatives) inherent in standard classification systems by identifying over 1 million environmentally relevant patents that were missed by conventional classifications (i.e. $\sim 2.9\%$ of the non-green patents sample). While our conservative text-based approach minimizes Type I errors through stringent contextual requirements and co-occurrence rules, the substantial expansion of green patent identification suggests that traditional methods systematically underestimate the scope of environmental innovation.